1. Find the length of a chorter leg of a right triangle if the longer leg is 10 miles more than the shorter leg and the hypotemuse is 10 miles less than twice the shorter leg.

2. The sum of two numbers is twenty-five and the sum of their square is three hundred and five. Find the number.

Answer by checkley75(3666)

X=SHORTER LEG
X+10=LONGER LEG
2X-10=HYPOTENUSE
X^2+(X+10)^2=(2X-10)^2
X^2+X^2+20X+100=4X^2-40X+100
2X^2-4X^2+60X=0
-2X^2+60X=0
2X^2-60x=0
x(2x-60)=0
x=0
2x-60=0
2x=60
x=60/2
x=30 answer.
proof
30^2+(30+10)^2=(2*30-10)^2
900+1600=(50)^2
2500=2500